# Thermodynamic topological classes of the rotating, accelerating black holes

### v1

@@ Abstract
In this paper, we extend our previous work [D. Wu, Phys. Rev. D 108, 084041 (2023)] to more general cases by including a rotation parameter. We investigate the topological numbers for the rotating, accelerating neutral black hole and its AdS extension, as well as the rotating, accelerating charged black hole and its AdS extension. We find that the topological number of an asymptotically flat accelerating black hole consistently differs by one from that of its non-accelerating counterpart. Furthermore, we show that for an asymptotically AdS accelerating black hole, the topological number is reduced by one compared to its non-accelerating AdS counterpart. In addition, we demonstrate that within the framework of general relativity, the acceleration parameter and the negative cosmological constant each independently add one to the topological number. However, when both factors are present, their effects neutralize each other, resulting in no overall change to the topological number.
Keywords:
Black Holes, Spacetime Singularities, AdS-CFT Correspondence
1 Introduction
@@
The study of the properties of black hole spacetimes and gravitational fields fundamentally relies on exact solutions to the gravitational field equations. Among the well-known family of exact black hole solutions to the four-dimensional Einstein field equations in General Relativity (GR), the most general Petrov type-D electrovacuum black hole solution is the Plebanski-Demianski metric AP98-98 . The simplest case of this metric, non-rotating, uncharged, and without NUT charge or cosmological constant, is referred to as the C-metric PRD2-1359 ; IJMPD15-335 . This class of solutions is also known as the accelerating black holes, which were interpreted by Kinnersley and Walker PRD2-1359 as the gravitational field produced by a point mass undergoing uniform acceleration, i.e., a pair of black holes accelerating away from each other in opposite directions GRG15-535 . In recent years, a wide range of studies has focused on the different features of accelerating black holes, including their global causal structure CQG23-6745 , quasinormal modes PRD109-084049 ; JHEP1022047 ; JHEP0224140 ; JHEP0224189 , quantum thermal properties PLA209-6 , holographic heat engines EPJC78-645 ; JHEP0219144 , weak/strong cosmic censorship conjecture PLB849-138433 ; SCPMA66-280412 , black hole shadows PRD103-025005 , and holographic complexity PLB823-136731 ; PLB838-137691 , etc JHEP0111114 ; JHEP1214057 ; JHEP0916082 ; PRL126-111601 ; JHEP0122102 ; PRL130-091603 ; JHEP0522063 ; JHEP1022074 ; JHEP0923122 ; JHEP1123073 ; JHEP0324050 ; JHEP0324079 ; 2404.04691 . Especially, in Refs. PRL117-131303 ; JHEP0517116 ; PRD98-104038 ; JHEP0419096 ; PLB796-191 ; CQG38-145031 ; CQG38-195024 ; PRD104-086005 ; 2306.16187 ; 2407.21329 , the thermodynamics of four-dimensional accelerating (charged and rotating) AdS black holes was thoroughly studied, leading to extensions of the first law of thermodynamics PRD7-2333 ; PRD13-191 , the Bekenstein-Smarr mass formula PRL30-71 , and the Christodoulou-Ruffini-like squared-mass formula PRL25-1596 ; PRD4-3552 ; PRD101-024057 ; PRD102-044007 ; PRD103-044014 ; JHEP1121031 to such spacetimes. [^2] Recently, these formulas were perfectly extended to the Lorentzian Taub-NUT spacetimes PRD100-101501 ; PRD105-124013 ; PRD108-064034 ; PRD108-064035 ; PLB846-138227 .
@@
The development of the above mass formulas represents just one facet of black hole thermodynamics. Quite recently, topology has emerged as a powerful mathematical tool, attracting substantial interest for its role in analyzing the thermodynamic phase transitions of black holes PRD105-104003 ; PRD105-104053 ; PLB835-137591 ; PRD107-046013 ; PRD107-106009 ; JHEP0623115 ; 2305.05595 ; 2305.05916 ; 2305.15674 ; 2305.15910 ; 2306.16117 ; PRD106-064059 ; PRD107-044026 ; PRD107-064015 ; 2212.04341 ; 2302.06201 ; 2304.14988 ; 2309.00224 ; 2312.12784 ; 2402.18791 ; 2403.14730 ; 2404.02526 ; 2407.09122 ; 2407.20016 ; 2408.03090 ; 2408.03126 ; NPB1006-116653 ; 2408.05870 . [^3] Topology has also proven effective in the study of the properties of light rings PRL119-251102 ; PRL124-181101 ; PRD102-064039 ; PRD103-104031 ; PRD104-044019 ; PRD105-024049 ; PRD105-064070 ; PRD108-104041 ; PRD109-064050 ; 2408.05569 and timelike circular orbits PRD107-064006 ; JCAP0723049 ; 2406.13270 . In Ref. PRL129-191101 , Wei et al. proposed a novel and systematic classification scheme for black holes by considering them as thermodynamic topological defects, categorized according to their respective topological numbers in a pioneering manner. A brief account of this method is presented below.
@@
where the extra variable $\tau$ can be interpreted as the inverse temperature of the cavity surrounding the black hole. The generalized Helmholtz free energy only exhibits on-shell characteristics and returns to the standard Helmholtz free energy of the black hole when $\tau=T^{-1}$. Now, using Duan’s $\phi$-mapping theory SS9-1072 ; NPB514-705 ; PRD61-045004 , one can construct a two-dimensional vector field $\phi=(\phi^{r},\phi^{\Theta})$ from the generalized Helmholtz free energy $\mathcal{F}$, where
@@
in which $r_{h}$ is the radius of the black hole’s event horizon, and $\Theta$ is an additional factor, with $\Theta\in[0,+\infty]$. It is important to note that the component $\phi^{\Theta}$ becomes infinite at $\Theta=0$ and $\Theta=\pi$, indicating that the vector points outward in both cases. Therefore, the unit vector is formulated as $n=(n^{r},n^{\Theta})$ with $n^{r}=\phi^{r_{h}}/||\phi||$ and $n^{\Theta}=\phi^{\Theta}/||\phi||$. The topological current $j^{\mu}$ satisfies the continuity equation $\partial_{\mu}j^{\mu}=0$, indicating that it is conserved. This current only has a nonzero contribution at points where the vector field $\phi^{a}(x_{i})$ equals zero, ensuring that the topological charge is well-defined. For a given parameter region $\rho$, the topological number is expressed as
@@
where $w_{i}$ represents the winding number at the $i$th point of $\phi$, $j^{0}$ denotes the density of the topological current, $\beta_{i}$ is the positive Hopf index, and the Brouwer degree $\eta_{i}=\mathrm{sign}(J^{0}({\phi}/{x})_{z_{i}})=\pm 1$, respectively. Note that the local winding number $w_{i}$ is regarded as a crucial tool for determining local thermodynamic stability. A positive $w_{i}$ suggests that the black holes are thermodynamically stable, while a negative $w_{i}$ indicates instability. The global topological number $W$ represents the difference between the counts of thermodynamically stable and unstable black holes at a specific temperature.
The elegance and general applicability of this method led to its widespread adoption, facilitating the exploration of topological numbers across different black hole models PRD107-064023 ; JHEP0123102 ; PRD107-024024 ; PRD107-084002 ; PRD107-084053 ; 2303.06814 ; 2303.13105 ; 2304.02889 ; 2306.13286 ; 2304.05695 ; 2306.05692 ; 2306.11212 ; EPJC83-365 ; 2306.02324 ; 2307.12873 ; 2309.14069 ; AP458-169486 ; 2310.09602 ; 2310.09907 ; 2310.15182 ; 2311.04050 ; 2311.11606 ; 2312.04325 ; 2312.06324 ; 2312.13577 ; 2312.12814 ; PS99-025003 ; 2401.16756 ; 2402.00106 ; PLB856-138919 ; AP463-169617 ; PDU44-101437 ; 2403.14167 ; PDU46-101617 ; 2405.02328 ; 2405.07525 ; 2405.20022 ; 2406.08793 ; AC48-100853 ; 2407.05325 ; 2408.08325 ; 2409.04997 ; EPJP139-806 . However, all previous research has primarily focused on non-accelerating cases, overlooking the topological numbers of accelerating black holes. Very recently, we have extended the topological approach to include non-rotating, accelerating black holes, considering both neutral and charged cases, as well as their AdS extensions PRD108-084041 . We have reached some intriguing conclusions: I) the topological number of an asymptotically flat, non-rotating, accelerating black hole always differs by one from that of its corresponding asymptotically flat, non-rotating, non-accelerating black hole; and II) the topological number of an asymptotically AdS, non-rotating, accelerating black hole consistently differs by minus one from that of its corresponding asymptotically AdS, non-rotating, non-accelerating black hole. Therefore, it makes sense to extend this work to more general rotating, accelerating black hole cases to investigate how the acceleration parameter affects the topological numbers of rotating black holes, and to verify whether these conclusions can be generalized to rotating, accelerating black holes. This also serves as the motivation for the current study.
@@
The remaining part of this paper is organized as follows. In section 2, we briefly introduce the rotating, accelerating charged AdS black hole. Section 3 begins with an exploration of the topological number of the rotating, accelerating black hole, focusing on the Kerr C-metric solution, and then extends this analysis to the Kerr-AdS C-metric solution, where a negative cosmological constant is present. In section 4, we examine the topological number of the rotating, accelerating charged black hole through the Kerr-Newman C-metric (KN C-metric) solution and subsequently extend this investigation to the KN-AdS C-metric case. Finally, section 5 presents our conclusions and future outlooks.
2 A brief introduction of rotating charged AdS accelerating black hole
@@
3 Topological classes of rotating neutral accelerating black hole
@@
3.1 Kerr C-metric black hole
@@
It is a simple matter to check that the above thermodynamic quantities simultaneously satisfy the first law and the Bekenstein-Smarr mass formula
@@
The Gibbs free energy for this black hole can be determined through the computation of the Euclidean action as outlined below
$\mathcal{I}_{E}=\frac{1}{16\pi}\int_{M}d^{4}x\sqrt{g}R+\frac{1}{8\pi}\int_{\partial{}M}d^{3}x\sqrt{h}(\mathcal{K}-\mathcal{K}_{0})\,.$ (15)
Here, $h$ represents the determinant of the induced metric $h_{ij}$, $\mathcal{K}$ denotes the extrinsic curvature of the boundary, and $\mathcal{K}_{0}$ corresponds to the extrinsic curvature of the massless C-metric solution, used as the reference background. The evaluation of the Euclidean action integral leads to the following expression for the Gibbs free energy
$G=\frac{\mathcal{I}_{E}}{\beta}=\frac{m(1-A^{2}a^{2})}{2\alpha K(1+A^{2}a^{2})}=\frac{M}{2}=M-TS-\Omega\,J\,,$ (16)
where $\beta=1/T$ being the interval of the time coordinate. Moreover, the final equality in Eq. (16) holds by applying the results from Eq. (3.1). Consequently, the conical singularity, characterized by the $(\lambda_{\pm}-\mu_{\pm})$-pairs, does not influence the Gibbs free energy calculation for the Kerr C-metric black hole.
In the following, we will derive the topological number of the Kerr C-metric black hole. We note that the Helmholtz free energy is simply given by
$F=G+\Omega\,J=M-TS\,.$ (17)
Replacing $T$ with $1/\tau$ in Eq. (17) and utilizing $m=(r_{h}^{2}+a^{2})/(2r_{h})$, then the generalized off-shell Helmholtz free energy is
@@
Figure 2: Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$ and $Ar_{0}=0.5$. There is one thermodynamically stable Kerr C-metric black hole for any value of $\tau$. Obviously, the topological number is: $W=1$.
Figure 3: The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the Kerr C-metric black hole with $a/r_{0}=1$, $Ar_{0}=0.5$, and $\tau/r_{0}=20$. The zero point (ZP) marked with black dot is at $(r_{h}/r_{0},\Theta)=(3.44,\pi/2)$. The blue contours $C$ is closed loop enclosing the ZP.
Assuming $Ar_{0}=0.5$, $a/r_{0}=1$ for the Kerr C-metric black hole, we plot Figs. 2 and 3 to illustrate key aspects of the system. These figures display the zero points of the component $\phi^{r_{h}}$ and the behavior of the unit vector field $n$ on a segment of the $\Theta-r_{h}$ plane, with $\tau=20r_{0}$, where $r_{0}$ is an arbitrary length scale defined by the size of the cavity enclosing the rotating, accelerating black hole.
@@
In Fig. 3, the zero point is found at $(r_{h}/r_{0},\Theta)=(3.44,\pi/2)$. As a result, the winding number $w_{i}$ for the blue contour $C$ is identified as $w_{1}=1$, which differ from those associated with the four-dimensional Kerr black hole PRD107-024024 . Regarding the global topological characteristics, the topological number $W=1$ for the Kerr C-metric black hole can be directly seen in Fig. 3, distinguishing it from the topological number of the Kerr black hole ($W=0$). Therefore, it can be inferred that the Kerr C-metric black hole and the Kerr black hole differ not only in geometric topology but also belong to distinct categories in thermodynamic topology. Consequently, it would be compelling to investigate the topological characteristics of black holes with unusual geometries, such as the type-D NUT C-metric black hole 2409.06733 , though this first requires establishing its consistent thermodynamics. In addition, combining the intriguing conclusion presented in Ref. PRD108-084041 , which states that the difference in topological numbers between an asymptotically flat, static, accelerating black hole and its corresponding non-accelerating counterpart is always unity, we suggest a more general conjecture within the framework of general relativity. Specifically, we propose that the difference in topological numbers between any asymptotically flat accelerating black hole and its corresponding non-accelerating black hole is consistently unity. This inference will be further substantiated in section 4.1, where we compute the topological number of the KN C-metric black hole and compare it with that of the KN black hole, thereby demonstrating the validity and applicability of this conjecture.
3.2 Kerr-AdS C-metric black hole

### v2

@@ Abstract
In this paper, we investigate the topological numbers for the rotating, accelerating neutral black hole and its AdS extension, as well as the rotating, accelerating charged black hole and its AdS extension. We find that the topological number of an asymptotically flat accelerating black hole consistently differs by one from that of its non-accelerating counterpart. Furthermore, we show that for an asymptotically AdS accelerating black hole, the topological number is reduced by one compared to its non-accelerating AdS counterpart. In addition, we demonstrate that within the framework of general relativity, the acceleration parameter and the negative cosmological constant each independently add one to the topological number. However, when both factors are present, their effects neutralize each other, resulting in no overall change to the topological number.
I Introduction
@@
The study of the properties of black hole spacetimes and gravitational fields fundamentally relies on exact solutions to the gravitational field equations. Among the well-known family of exact black hole solutions to the four-dimensional Einstein field equations in General Relativity (GR), the most general Petrov type-D electrovacuum black hole solution is the Plebanski-Demianski metric AP98-98 . The simplest case of this metric, non-rotating, uncharged, and without NUT charge or cosmological constant, is referred to as the C-metric PRD2-1359 ; IJMPD15-335 . This class of solutions is also known as the accelerating black holes, which were interpreted by Kinnersley and Walker PRD2-1359 as the gravitational field produced by a point mass undergoing uniform acceleration, i.e., a pair of black holes accelerating away from each other in opposite directions GRG15-535 . In recent years, a wide range of studies has focused on the different features of accelerating black holes, including their global causal structure CQG23-6745 , quasinormal modes PRD109-084049 ; JHEP1022047 ; JHEP0224140 ; JHEP0224189 ; 2305.04040 ; 2408.05957 , quantum thermal properties PLA209-6 , holographic heat engines EPJC78-645 ; JHEP0219144 , weak/strong cosmic censorship conjecture PLB849-138433 ; SCPMA66-280412 , black hole shadows PRD103-025005 , and holographic complexity PLB823-136731 ; PLB838-137691 , etc PRL126-111601 ; JHEP0122102 ; PRL130-091603 ; JHEP1022074 ; JHEP0324050 ; JHEP0324079 ; 2404.04691 ; 1711.04522 ; 1906.09478 ; 2403.09757 ; 1408.3124 ; 2205.15777 ; 2211.06517 ; 2305.03765 ; 2309.13656 ; 2411.12513 . Especially, in Refs. PRL117-131303 ; JHEP0517116 ; PRD98-104038 ; JHEP0419096 ; PLB796-191 ; CQG38-145031 ; CQG38-195024 ; PRD104-086005 ; 2306.16187 ; 2407.21329 ; 2406.01897 , the thermodynamics of four-dimensional accelerating (charged and rotating) AdS black holes was thoroughly studied, leading to extensions of the first law of thermodynamics PRD7-2333 ; PRD13-191 , the Bekenstein-Smarr mass formula PRL30-71 , and the Christodoulou-Ruffini-like squared-mass formula PRL25-1596 ; PRD4-3552 ; PRD101-024057 ; PRD102-044007 ; PRD103-044014 ; JHEP1121031 to such spacetimes. [^1] Recently, these formulas were perfectly extended to the Lorentzian Taub-NUT spacetimes PRD100-101501 ; PRD105-124013 ; PRD108-064034 ; PRD108-064035 ; PLB846-138227 .
@@
The development of the above mass formulas represents just one facet of black hole thermodynamics. Quite recently, topology has emerged as a powerful mathematical tool, attracting substantial interest for its role in analyzing the thermodynamic phase transitions of black holes PRD105-104003 ; PRD105-104053 ; PLB835-137591 ; PRD107-046013 ; PRD107-106009 ; JHEP0623115 ; 2305.05595 ; 2305.05916 ; 2305.15910 ; PRD106-064059 ; PRD107-044026 ; PRD107-064015 ; 2212.04341 ; 2302.06201 ; 2304.14988 ; 2309.00224 ; 2312.12784 ; 2402.18791 ; 2404.02526 ; 2407.09122 ; 2407.20016 .  [^2] Topology has also proven effective in the study of the properties of light rings PRL119-251102 ; PRL124-181101 ; PRD102-064039 ; PRD103-104031 ; PRD104-044019 ; PRD105-024049 ; PRD105-064070 ; PRD108-104041 ; PRD109-064050 ; 2408.05569 ; 2412.18083 and timelike circular orbits PRD107-064006 ; JCAP0723049 ; 2406.13270 . In Ref. PRL129-191101 , Wei et al. proposed a novel and systematic classification scheme for black holes by considering them as thermodynamic topological defects, categorized according to their respective topological numbers in a pioneering manner. A brief account of this method is presented below.
@@
where the extra variable $\tau$ can be interpreted as the inverse temperature of the cavity surrounding the black hole.  [^3] In a recent preprint 2504.00503 , the authors discovered two special cases of the generalized Helmholtz free energy, corresponding to the generalized Helmholtz free energy in bumblebee gravity PRD39-683 and three-dimensional new massive gravity PRL102-201301 ; RMP84-671 . Therefore, when addressing the thermodynamic topology classification of black holes in these two theories, the form of the generalized free energy may require modification. The generalized Helmholtz free energy only exhibits on-shell characteristics and returns to the standard Helmholtz free energy of the black hole when $\tau=T^{-1}$. According to Ref. PRL129-191101 , one can construct a two-dimensional vector field $\phi=(\phi^{r},\phi^{\Theta})$ from the generalized Helmholtz free energy $\mathcal{F}$, where
@@
in which $r_{h}$ is the radius of the black hole’s event horizon, and $\Theta$ is an additional factor, with $\Theta\in[0,+\infty]$. It is important to note that the component $\phi^{\Theta}$ becomes infinite at $\Theta=0$ and $\Theta=\pi$, indicating that the vector points outward in both cases.
Using Duan’s $\phi$-mapping topological current theory SS9-1072 ; NPB514-705 ; PRD61-045004 , a topological current can be described as follows  [^4] Unlike topological methods based on light rings PRL119-251102 ; PRL124-181101 or timelike circular orbits PRD107-064006 , which prioritize dynamical or geometric properties, the $\phi$-mapping theory offers distinct advantages in probing thermodynamic features. Specifically, our work employs $\phi$-mapping theory to classify black hole solutions by analyzing the global topology of thermodynamic defects, establishing a unified framework for black hole solution classification PRL129-191101 . This contrasts with other applications of $\phi$-mapping theory that focus on characterizing critical points during black hole phase transitions PRD105-104003 . While these alternative approaches excel at dissecting phase behavior, our work emphasizes the intrinsic classification of black hole solutions through their thermodynamic defect structures, thereby complementing existing analyses of dynamic or geometric traits.:
$j^{\mu}=\frac{1}{2\pi}\epsilon^{\mu\nu\rho}\epsilon_{ab}\partial_{\nu}n^{a}\partial_{\rho}n^{b}\,,\qquad\mu,\nu,\rho=0,1,2,$ (3)
where $\partial_{\nu}=\partial/\partial{}x^{\nu}$ and $x^{\nu}=(\tau,~{}r_{h},~{}\Theta)$. The unit vector $n$ formulates as $n=(n^{r},n^{\Theta})$, where $n^{r}=\phi^{r_{h}}/||\phi||$ and $n^{\Theta}=\phi^{\Theta}/||\phi||$. It is easy to indicate that the topological current (3) given above is conserved, allowing one to easily deduce $\partial_{\mu}j^{\mu}=0$. It is then demonstrate that the topological current $j^{\mu}$ is a $\delta$-function of the field configuration NPB514-705 ; PRD61-045004 :
$j^{\mu}=\delta^{2}(\phi)J^{\mu}\Big{(}\frac{\phi}{x}\Big{)}\,,$ (4)
where the 3-dimensional Jacobian $J^{\mu}\big{(}\phi/x\big{)}$ is defined as: $\epsilon^{ab}J^{\mu}\big{(}\phi/x\big{)}=\epsilon^{\mu\nu\rho}\partial_{\nu}\phi^{a}\partial_{\rho}\phi^{b}$. It is easy to obtain that $j^{\mu}$ equals to zero only when $\phi^{a}(x_{i})=0$. Therefore, for a given parameter region $\rho$, the topological number is expressed as
@@
where $w_{i}$ represents the winding number at the $i$th point of $\phi$, $j^{0}$ denotes the density of the topological current, $\beta_{i}$ is the positive Hopf index, and the Brouwer degree $\eta_{i}=\mathrm{sign}(J^{0}({\phi}/{x})_{z_{i}})=\pm 1$, respectively. Note that the local winding number $w_{i}$ is regarded as a crucial tool for determining local thermodynamic stability. The positive or negative heat capacity of a black hole state can be shown to, respectively, correspond to a positive or negative winding number, $w=+1$ and $w=-1$, associated with locally stable and unstable black hole states. The global topological number $W$ represents the difference between the counts of thermodynamically stable and unstable black holes at a specific temperature. Therefore, the global topological number $W$ characterizes each black hole system and can be used as a classification parameter.
The elegance and general applicability of this method led to its widespread adoption, facilitating the exploration of topological numbers across different black hole models PRD107-064023 ; JHEP0123102 ; PRD107-024024 ; PRD107-084002 ; PRD107-084053 ; 2303.06814 ; 2304.05695 ; 2306.05692 ; 2306.11212 ; EPJC83-365 ; 2306.02324 ; AP458-169486 ; 2310.09602 ; 2310.15182 ; 2311.04050 ; 2311.11606 ; 2312.04325 ; 2312.06324 ; 2312.13577 ; 2312.12814 ; PS99-025003 ; 2402.00106 ; PLB856-138919 ; AP463-169617 ; PDU44-101437 ; 2403.14167 ; PDU46-101617 ; 2405.02328 ; 2405.07525 ; 2405.20022 ; 2406.08793 ; 2407.05325 ; 2408.08325 ; 2409.04997 . However, all previous research has primarily focused on non-accelerating cases, overlooking the topological numbers of accelerating black holes. Very recently, we have extended the topological approach to include non-rotating, accelerating black holes, considering both neutral and charged cases, as well as their AdS extensions PRD108-084041 . Extending this work to more general rotating, accelerating black holes will clarify how the acceleration parameter influences their topological numbers. This also serves as the motivation for the current study.
@@
Our analysis now emphasizes that prior studies focused on non-accelerating black holes or static accelerating cases, while this work extends the topological classification to rotating, accelerating black holes in both asymptotically flat and AdS spacetimes. We also highlight the discovery that the acceleration parameter and cosmological constant counterbalance each other’s topological contributions–a novel result not explored in earlier work.
The remaining part of this paper is organized as follows. In section II, we briefly introduce the rotating, accelerating charged AdS black hole. Section III begins with an exploration of the topological number of the rotating, accelerating black hole, focusing on the Kerr C-metric solution, and then extends this analysis to the Kerr-AdS C-metric solution, where a negative cosmological constant is present. In section IV, we examine the topological number of the rotating, accelerating charged black hole through the Kerr-Newman C-metric (KN C-metric) solution and subsequently extend this investigation to the KN-AdS C-metric case. Finally, section V presents our conclusions and future outlooks.
II A brief introduction of rotating charged AdS accelerating black hole
@@
III Topological classes of rotating neutral accelerating black hole
@@
III.1 Kerr C-metric black hole
@@
Here and hereafter (in Eqs. (LABEL:ThermKerrAdSC), (LABEL:ThermKNC), and (LABEL:ThermKNAdSC)), $M$, $S$, $T$, $\Omega$, and $J$ represent the mass, Bekenstein-Hawking entropy, Hawking temperature, angular velocity, and angular momentum, respectively; $\mu_{\pm}$ are the tensions of the conical deficits on the north and south poles, and $\lambda_{\pm}$ are the thermodynamic lengths conjugate to the tensions $\mu_{\pm}$. It is a simple matter to check that the above thermodynamic quantities simultaneously satisfy the first law and the Bekenstein-Smarr mass formula
@@
In the following, we will derive the topological number of the Kerr C-metric black hole. The generalized off-shell Helmholtz free energy is
@@
Figure 2: (a) Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$ and $Ar_{0}=0.5$. There is always one thermodynamically stable four-dimensional Kerr C-metric black hole for any value of $\tau$. Obviously, the topological number is: $W=1$; (b) The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the four-dimensional Kerr C-metric black hole with $a/r_{0}=1$, $Ar_{0}=0.5$, and $\tau/r_{0}=20$. The zero point (ZP) marked with black dot is at $(r_{h}/r_{0},\Theta)=(3.44,\pi/2)$. The blue contours $C$ is closed loop enclosing the ZP.
Assuming $Ar_{0}=0.5$, $a/r_{0}=1$ for the Kerr C-metric black hole, we plot Figs. 2 and 2 to illustrate key aspects of the system. These figures display the zero points of the component $\phi^{r_{h}}$ and the behavior of the unit vector field $n$ on a segment of the $\Theta-r_{h}$ plane, with $\tau=20r_{0}$, where $r_{0}$ is an arbitrary length scale defined by the size of the cavity enclosing the rotating, accelerating black hole.
@@
In Fig. 2, the zero point is found at $(r_{h}/r_{0},\Theta)=(3.44,\pi/2)$. As a result, the winding number $w_{i}$ for the blue contour $C$ is identified as $w_{1}=1$, which differ from those associated with the four-dimensional Kerr black hole PRD107-024024 . Regarding the global topological characteristics, the topological number $W=1$ for the Kerr C-metric black hole can be directly seen in Fig. 2, distinguishing it from the topological number of the Kerr black hole ($W=0$). Therefore, it can be inferred that the Kerr C-metric black hole and the Kerr black hole differ not only in geometric topology but also belong to distinct categories in thermodynamic topology. The underlying mechanism responsible for these distinct properties may be attributed to the acceleration parameter, which introduces additional geometric effects via conical deficits. This modifies the root structure of $f(r)$, consequently reducing the number of unstable branches. This physical intuition suggests that acceleration could suppress the complexity of thermodynamic phase transitions by stretching spacetime along the symmetry axis. Consequently, it would be compelling to investigate the topological characteristics of black holes with unusual geometries, such as the type-D NUT C-metric black hole 2409.06733 , though this first requires establishing its consistent thermodynamics. In addition, combining the intriguing conclusion presented in Ref. PRD108-084041 , which states that the difference in topological numbers between an asymptotically flat, static, accelerating black hole and its corresponding non-accelerating counterpart is always unity, we suggest a more general conjecture within the framework of general relativity. Specifically, we propose that the difference in topological numbers between any asymptotically flat accelerating black hole and its corresponding non-accelerating black hole is consistently unity. This inference will be further substantiated in section IV.1, where we compute the topological number of the KN C-metric black hole and compare it with that of the KN black hole, thereby demonstrating the validity and applicability of this conjecture.
III.2 Kerr-AdS C-metric black hole